OpenStudy (anonymous):

Solve Z = 9x + 9y for y. please help:)

OpenStudy (anonymous):

Try moving 9x on the left side and then dividing by 9. See where does that get you.

OpenStudy (anonymous):

move or subtract ?

OpenStudy (suwhitney):

Subtract the 9x from the right side to left side.

OpenStudy (cgreenwade2000):

Moving never means just moving it. With equations you have to balance out both sides to keep it all equal. So since you are adding 9x on both sides now, you want to subtract it on both sides.

OpenStudy (suwhitney):

Then move the Z to the right side. So now you should have -9x=-z+9y

OpenStudy (anonymous):

\[ z = 9x + 9y\\ z-9x=9y\\ y=\frac{z-9x}{9} \]

OpenStudy (anonymous):

ooh... i feel dumb

OpenStudy (anonymous):

Thanks